OpenStudy (anonymous):

Please help.. What are the solutions to the following system of equations? y = x2 + 12x + 30 8x − y = 10 (−4, −2) and (2, 5) (−2, −4) and (2, 5) (−2, −4) and (5, 2) No Real Solutions

OpenStudy (anonymous):

@Nurali

OpenStudy (uri):

@DebbieG Help her ^_^ Uri is proud of you.

OpenStudy (debbieg):

What methods have you learned for solving a system? Substitution, elimination?

OpenStudy (anonymous):

Elimination I think

OpenStudy (anonymous):

And some substitution but I cant remember how to do that method

OpenStudy (debbieg):

I would probably opt for substitution here, since you have a quadratic (a parabola) and a linear equation (a line). But either would work. If you solve the 2nd equation for y, then you have two equations that both give y as a function of x, right? E.g., you have y={some stuff} and y={some different stuff} so since y= both of those things, you can set them equal to each other.

OpenStudy (debbieg):

When you do that, you'll have a single, quadratic equation to solve.

OpenStudy (anonymous):

..... that makes no sense

OpenStudy (debbieg):

What part didn't make sense? Here are your two equations: \[y=x^2+12x+30\]\[8x-y=10\]Can you solve the 2nd one for y?

OpenStudy (anonymous):

Idk how.. I only know how to solve the ones with Y on the other side of the equal sign

OpenStudy (debbieg):

You just need to rearrange 8x-y=10 so that the y is by itself. Try adding y to both sides: 8x-y+y=10+y What do you get when you simplify that? Then subtract 10 from both sides.

OpenStudy (anonymous):

IM SO FREAKING LOST

OpenStudy (anonymous):

THANKS GUYS. I don't need any more help

OpenStudy (debbieg):

@jessssloukaa the idea is not to do the problem FOR her.